OpenStudy (owlcoffee):

Find the equation of tangent lines that go through point P((3,-5) to the circumference x^2+y^2-x+3y-2=0

OpenStudy (anonymous):

Find the derivative using implicit differentiation \[ 2 x + 2 y y' -1 + 3 y'=0\\ 2 y y' + 3 y'= 1-2 x \\ y'( 2y +3) = 1-2 x \\ y'= \frac { 1-2 x}{( 2y +3)} \]

OpenStudy (anonymous):

Find y' at the point P((3,-5) \[ y'= \frac {1 -6} {-10+3}=\frac{-5} {-7}=\frac 5 7 \]

OpenStudy (anonymous):

Now use \[ y-y_0 = \frac 5 7 (x-x0) \]

OpenStudy (anonymous):

At (3,-5) \[y-y_0 = \frac 5 7 (x-x0)\\ y+5 =\frac 5 7 (x-3)\\ y =-5+\frac 5 7 (x-3) \]

OpenStudy (anonymous):

Did you get it?

OpenStudy (owlcoffee):

yeah, I got curious about that implicit differentiation. What does it mean geometrically?

OpenStudy (anonymous):

To be able to find the equation of the tangent, you have to find the derivative.